OpenStudy (anonymous):

Solve for x: x^2 + 24x + 90 = 0

OpenStudy (saifoo.khan):

Use the quadratic formula.

OpenStudy (anonymous):

I don't understand the quadratic formula, the lessons aren't very helpful.

OpenStudy (saifoo.khan):

Use, \[x = \frac{-b \pm \sqrt{b^2-4ac }}{2a}\]

OpenStudy (saifoo.khan):

The value of a = 1 b = 24 c = 90/

OpenStudy (saifoo.khan):

INsert the values in their places to get the final numerical values.

OpenStudy (jamesj):

To see an example, watch this slightly goofy guy: http://www.youtube.com/watch?v=s80J2dAUUyI

OpenStudy (anonymous):

Oh okay hold on..

OpenStudy (anonymous):

a is the coefficient of the x^2 value. b is the coefficient of the x value. c is the constant.

OpenStudy (anonymous):

I got a strange answer. I don't know if it's right.

OpenStudy (jamesj):

the answer will be "strange" in as much as it's not an integer.

OpenStudy (anonymous):

Well i got \[-72 \pm \sqrt{3}\] Is that right or no?..

OpenStudy (jamesj):

No, that can't be right because the -72 just has nothing to do with -b/2 where b = 24. And in fact it isn't right. Calculate it again.

OpenStudy (anonymous):

No because the original equation i received was \[-12 \pm \sqrt{108}\]

OpenStudy (jamesj):

Still no. First, what is b^2 - 4ac for a = 1 b = 24 c = 90

OpenStudy (anonymous):

24^2 - 4(1)(90)

OpenStudy (jamesj):

\[b^2 - 4ac = 24^2 - 4(90) = 216 \]

OpenStudy (anonymous):

Oh I divided before i subtracted.

OpenStudy (jamesj):

So what's your answer now?

OpenStudy (anonymous):

\[-12 \pm \sqrt{18}\]

OpenStudy (jamesj):

the -12 is right. the sqrt(18) isn't

OpenStudy (anonymous):

Okay so then i'm not getting this.

OpenStudy (jamesj):

\[ \frac{\sqrt{b^2 - 4ac}{2a} = \frac{\sqrt{216}}{2} = \sqrt{216/4} = \sqrt{54} = 3\sqrt{6} \]

OpenStudy (jamesj):

hmm ... look like eqn editor is broken.

OpenStudy (anonymous):

Haha yeah..

OpenStudy (jamesj):

sqrt(b^2 - 4ac)/2a = sqrt(216)/2 = sqrt(216/4) = sqrt(54) = 3 sqrt(6)

OpenStudy (anonymous):

Why sqrt(216/4) ??

OpenStudy (jamesj):

because (1/x)sqrt(y) = sqrt(y/x^2)

OpenStudy (anonymous):

Oh okay so then the answer is 3 plus or negative √6 ?..

OpenStudy (jamesj):

No. I hate giving the answer, but I'm going to give it and ask you to make sure you can get it as well: \[-12 \pm 3 \sqrt{6}\]

OpenStudy (anonymous):

Ohh! So i was right but it just needed to be simplified from √18 to 3√6?

OpenStudy (jamesj):

\[\sqrt{18} \ne 3\sqrt{6}\]

OpenStudy (jamesj):

so you were not right.

OpenStudy (anonymous):

Okay never mind then.

OpenStudy (anonymous):

just use your damn scientific calculator.

OpenStudy (jamesj):

just use your brain actually because this isn't hard with some practice.

OpenStudy (aravindg):

hey jamess

OpenStudy (aravindg):

help me